OpenStudy (anonymous):

identify the restricted values a+2 _______ 2a-8

OpenStudy (unklerhaukus):

what values for \(a\) would mean the fraction is undefined

OpenStudy (anonymous):

any values that make the denominator equal to zero

Parth (parthkohli):

There's a good trick. Equate the denominator to zero!\[\rm 2a - 8 = 0\]Solve that equation to get your restricted value(s).

OpenStudy (unklerhaukus):

right @love1904 , can you find suce an \(a\)? use Parthy's trick

OpenStudy (anonymous):

yeah, I am trying

OpenStudy (anonymous):

a=4

Parth (parthkohli):

Very good!

OpenStudy (unklerhaukus):

is there any other value for \(a\) that would make the fraction undetermined ?

OpenStudy (anonymous):

-2

OpenStudy (unklerhaukus):

lets see what would happen if \(a=-2\) \[\frac{a+2}{2a-8}=\frac{-2+2}{2(-2)-8}=\frac{0}{-12}=\] is this fraction defined?

OpenStudy (anonymous):

no

Parth (parthkohli):

@love1904 Only the denominator—not numerator.

OpenStudy (unklerhaukus):

when \(a=-2\)\[\frac{a+2}{2a-8}=\frac{0}{-12}=0\] the fraction equals zero ( it is defined )

OpenStudy (anonymous):

Great!! Good Jobs Guys thanks.

OpenStudy (unklerhaukus):

so there is only one restricted value for a

OpenStudy (anonymous):

another one is coming

OpenStudy (anonymous):

Simplify 2t-12 __________________ 6-t

OpenStudy (unklerhaukus):

factorise the numerator and then cancel